You are here: Foswiki>FLAO Web>LbtLab>LabLog20100216 (17 Feb 2010, EnricoPinna)Edit Attach
POWER SHOCK last night AG electronic demaged

Electronic RE-CHECK

- Poweron & ping devices (ts8 39&47, bcu39&47) : OK
- Box fans: OK
- LittleJoe fans: can not test
- Temperature reading (internal & water): OK
- Box47 power supply:OK
- Bcu39 bank switch: cannot test
- Bcu39 reset: OK
- Bcu47 reset: OK
- Bcu47 bank switch: cannot test
- Flowerpot PIC: OK
- Flowerpot sensirion reading: OK (board not present)
- Cube rotator: OK
- Cube stage: OK
- lamp: OK
- Stages XYZ: all OK
- LJ 47: OK - frames live & quadrants (FRAME With SNOW pattern)
- LJ 39: OK - frames live & quadrants

BCU39 PROBLEM
ping stops and restart after restarting, but only the first time, than is not responding
Restarting again all the system the BCU is not responding at all

BOX39 opened, now BCU39 properly working
Also the snow-pattern on CCD47 disappear

Afternoon

W board mounted on the stages and connected
W CHECK
- FW1&2: OK
- ADC 1&2: OK
- Rerot: OK
- TTMirror: OK
- LC traslator: X -> OK, Y->??????

LIGHT source images on CCD47
A strange pattern on CCD47 (fingerprint on CCD window?)

We can not check light CCD39 because the BCU39 is not working

-- EnricoPinna - 16 Feb 2010
Topic revision: r3 - 17 Feb 2010, EnricoPinna
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback